Cheryl invests her inheritance in 2 separate accounts. She puts part of it in an account
earning 18% annual interest, and twice that amount in an account earning 12% interest. If
the account at 12% earns $1200 more interest for the year than the 18% account, what is
the amount of the inheritance?

Answer by Theo(13342)

x = amount invested at 18%
y = amount invested at 12%
.12 * y = .18 * x + 1200
amount invested at 12% is two times the amount invested at 18%, getting you
y = 2 * x, therefore:
.12 * 2 * x = .18 * x + 1200
simplify to get:
.24 * x = .18 * x + 1200
subtract .18 * x from both sides to get:
.06 * x = 1200
solve for x to get x = 20,000
to confirm:
20,000 invested at 18% gives you 3600 interest.
40,000 invested at 12% gives you 4800 interest.
4800 is 1200 more than 3600, confirming that the amount invested at 12% gives you 1200 more interest than the amount invested at 18%.